OpenStudy (anonymous):

A music store buys compact discs for $12 each and sells them for $15 each. What is the percent of markup on these compact discs?

OpenStudy (kropot72):

The markup on the cost price is \[\frac{15-12}{12}\times100=?\ percent\]

OpenStudy (kropot72):

You subtract the cost price ($12) from the selling price ($15) as follows: 15 - 12 = ? ..............(1) Then you divide the result of (1) by the cost price ($12) \[\frac{?}{12}=?\ ?............(2)\] Then multiply the result of (2) by 100 to find the percentage.